OpenStudy (anonymous):

3(3x-1)^2=21 solve the equation algebraically

OpenStudy (anonymous):

This is actually 3(3x -1)(3x-1) = 21 All you need to do is expand the brackets then simplify.

OpenStudy (anonymous):

x = 1.21525043702153

OpenStudy (anonymous):

Make sure you then shift the 21 so the right side equating the equation to 0 to solve.

OpenStudy (anonymous):

Can you do the rest?

OpenStudy (anonymous):

my answer sheet said that the correct answer ia \[\sqrt{7}/3 \pm1/3\]

OpenStudy (anonymous):

you can start be dividing both sides of the equation by 3 this will give you a new equations (3x-1)^2 = 7 then square root both sides 3x-1 = 7^(1/2) [or root 7] and 1 to both sides the divide by 3 gives you: x= ((sqr.root 7) + 1)/3 which is x = 1.21525043702153

OpenStudy (anonymous):

oh ok'